Donnie Hilfiger has two classes of stock authorized: $1 par preferred and $0.01 par value common. As of the beginning of 2021, 310 shares of preferred stock and 4,100 shares of common stock have been issued. The following transactions affect stockholders’ equity during 2021: March 1 Issue 1,200 shares of common stock for $43 per share. May 15 Purchase 410 shares of treasury stock for $36 per share. July 10 Resell 210 shares of treasury stock purchased on May 15 for $41 per share. October 15 Issue 210 shares of preferred stock for $46 per share. December 1 Declare a cash dividend on both common and preferred stock of $0.60 per share to all stockholders of record on December 15. (Hint: Dividends are not paid on treasury stock.) December 31 Pay the cash dividends declared on December 1. Donnie Hilfiger has the following beginning balances in its stockholders’ equity accounts on January 1, 2021: Preferred Stock, $310; Common Stock, $41; Additional Paid-in Capital, $76,500; and Retained Earnings, $30,700. Net income for the year ended December 31, 2021, is $11,000.
Answer #1

The stockholders' Equity section of the company for the year ended Dec 31, 2021 is shown as follows:-

Balance Sheet
(Stockholders' Equity Section)
December 31, 2021
Stockholders' Equity:
Preferred Stock (See note below) 520
Common Stock 53
Additional Paid-in Capital 138,588
Total Paid-in Capital 139,161
Retained Earnings 38,328
Treasury Stock (7,200)
Total Stockholders' Equity 170,289

Notes:-

1) The calculation of each item of stockholders' Equity is shown as follows:- (Amounts in $)

Preferred Stock Common Stock Additional Paid-in Capital Retained Earnings Treasury Stock
Beg. Balance, Jan 1, 2021 310 41 76,500 30,700 0
Issued common shares 0 12 (1,200*$0.01 par) 51,588 [1,200*(43-0.01)] 0 0
Purchase treasury stock 0 0 0 0 (14,760)(410*36)
Sale of Treasury Stock 0 0 1,050 [210*(41-36)] 0 7,560 (210*$36)
Issued preferred shares 210 (210*$1 par) 0 9,450 [210*(46-1)] 0 0
Cash dividends paid 0 0 0 (3,372) (see note 2)
Net Income 0 0 0 11,000 0
Ending Balance, Dec 31, 2021 520 53 138,588 38,328 (7,200)

2) Common share outstanding at the end

= Beginning+New Issue-Treasury stock purchased+Treasury stock resold

= 4,100+1,200-410+210 = 5,100 shares

Preferred shares outstanding at the end = Beginning+New issue

= 310+210 = 520

Dividend paid = (5,100+520)*$0.60 per share = $3,372